Tom had 67 points on his last math quiz, and it increased his average from 27 to 32 points. How many points on the last quiz would have increased his average to 35 points per quiz.
Answer by Boreal(15235)

average = total/n
27= total/n
after last quiz, his average is (total+67)/(n+1)=32
Therefore, 32 n+ 32=total +67
and 27n=total
So 32 n + 32=27 n +67
5n=35
n=7
He had 189 points in his first 7 quizzes.
after getting a 67, he had 256 points in 8 quizzes, which is 32 average.
If he had gotten a total of 280 points in 8 quizzes, he would have had a 35 average.
He would have needed a 280-189 score or a 91 on his last quiz.